The first divers used metal ‘diving bells’ to move air under the water surface.

(a) Modelling the air in bell as a hemisphere with diameter of 2.5 metres, use the following information to calculate the upthrust produced due to the water.

Ρwater= 997 kgm-3

Volume of a sphere = 4/3𝜋r3

g = 9.81 ms-2

(b) If the bell was fully submerged and then its suspension cable was released, what would its maximum acceleration be, given that its mass was 4500 kg?


1
Expert's answer
2022-03-28T16:13:42-0400

a)

"F=\rho gV=\rho g\frac{\pi d^3}{12}=40~kN,"

b)

"ma=mg-F,"

"a=g-\frac Fm=0.92~\frac m{s^2}."